Arne fixed the chiller on the calorimeter during yesterday's controlled access and will commission it this morning during the 31 MHz running.
At ~1030 we plan to request a controlled access to repair the broken harp scanner, swap out an ADC in the left arm, and the tech's were going to look into the intermittant communication issue with the left quads.
The schedule has regular CW production beam being restored to the Halls starting 1200. It will probably be closer to 1300 if we're lucky.
Once we come back up we'll do a Bull's eye scan and continue with the
production run plan.